OpenStudy (anonymous):

solve the following system and show all of the work x=2y=14 x+3y=9

razor99 (razor99):

thats simmultaneous equation

razor99 (razor99):

right?

OpenStudy (anonymous):

I don't no I came up with x=13, y=-1 but I don't no if that is right

razor99 (razor99):

cld u show the working how did u do it please

razor99 (razor99):

i want the second part

OpenStudy (anonymous):

would you recheck your post, Jean ! x=2y=14 x+3y=9

OpenStudy (anonymous):

I assume x - 2y = 14 then, y = -5 => x = 14 + 2(-5) = 24

OpenStudy (anonymous):

solution ( 24, -5)

OpenStudy (anonymous):

ok I'm doing something wrong can't figure it out